est Review 12 of A fea jumps st leaves the groungh up to a maxmum height of 0 370 m of free fa your answer in Hint(s) B Complete
Answer #1

Part A

By using Kinematic equation,

v^2 = u^2 + 2as

v^2 = 2gh

v = sqrt(2*9.8*0.370)

= 2.69 m/s

Answer